Suppose a firm produces 15 units per month and it has a total variable cost of 15000 per month. If its average fixed costs are 350 per month, what are its total costs per month
1
Expert's answer
2019-09-27T09:12:46-0400

TC = FC + VC = AFC×Q + VC = 350×15 + 15,000 = 20,250.